Calculate the density, specific weight, and volume of chloride gas at 25 °C and pressure of 600 000 N/m2 absolute.




1
Expert's answer
2022-02-22T15:37:01-0500

Given the pressure of chloride gas at 25 ° C is

P=600000N/m2

and 1 pascal is equal to 1 N/m2.

therefore P= 600000 pa


Convert 25 ° C to k 

T=273+25 =298K

Gas constant for chloride assume to be 117 kj/kg. K

take g to be g=9.81m/s2

(a) Density of chloride gas

From an ideal gas law P=ρRT

yo have P, R, T

density "\u03c1 =P\/RT"

ρ = 600000pa/(117 x103x298)

density=0.0172 kg/m3


(b) specific weight

weight=density x gravitational constant 

w= ρ g

= 0.0172 X 9.81

w=0.1687 N/m3

(c) Volume

V=I/ ρ 

V=1/0.0172

V=58.1395 m3/kg